You have 3/4 of a book to read. You need to finish it in 10 days.
a. What fraction of the book do you need to read each day?
b. The book has 120 total pages. How many pages do you need to read
each day?

Answer by ReadingBoosters(3246)

x = 120 pg book
3/4 over 10 days
(3/4) / 10, which is the same (3/4)(1/10)= 3/40 of the book to read daily
(3/4)x = (3/4)(120) = (3/1)(30) = 90 pages to read
90 pages / 10 days = 9 pages per day